OpenStudy (bluesea24):

how to solve: integation root of (3x+4)^3 dx.pls help...

OpenStudy (bluesea24):

this is the Q:\[\int\limits \sqrt{(3x + 4)^3 dx}\]

OpenStudy (anonymous):

make u substitution , for u=3x+4

OpenStudy (bluesea24):

so now eq will be \[\frac{ 1 }{ 3 }\ \int\limits \sqrt{u^3du}\]

OpenStudy (bluesea24):

what is final ans?

OpenStudy (anonymous):

yeah that would be the integral now you can rewrite that as \[\frac{ 1 }{3 }\int\limits_{}^{}u^{\frac{ 3 }{ 2 }}du\]

OpenStudy (bluesea24):

cool wow u did it.... thannnnnk uuuuuuu. so much :)

OpenStudy (anonymous):

no problem ur welcome :)

OpenStudy (bluesea24):

i got ans as \[1/4*\log u ^{5\div6}+c\]

OpenStudy (bluesea24):

Q is :\[\int\limits dx \div \sqrt[3]{(4x-3)^\frac{ 5 }{ 2 }}\]

OpenStudy (bluesea24):

i got above ans for this Q. ans is right?

OpenStudy (anonymous):

give me a min to check

OpenStudy (anonymous):

you got mistake \[\int\limits_{}^{}{\frac{ du }{ u^{5/6} }}\] is not \[\log{u^{\frac{ 5 }{ 6 }}}\]

OpenStudy (anonymous):

when you get to that integral you rewrite it as \[\int\limits_{}^{}u^{-\frac{ 5 }{ 6 }}du\] and then you solve it with the formula \[\int\limits_{}^{}x^{n}dx=\frac{ x^{n+1} }{ n+1 }\]

OpenStudy (bluesea24):

ok got it.... silly mistake there. thanks once again...

OpenStudy (anonymous):

no problem ^^

OpenStudy (bluesea24):

why is \[\int\limits \frac{ dx }{ 1+x^2 }=\tan^{-1} x=-\cos^{-1}x\]

OpenStudy (bluesea24):

how tan−1x=−cos−1x in above formula?

OpenStudy (anonymous):

hm i am not sure if that is correct , never came across it

OpenStudy (bluesea24):

Q is:\[\int\limits \frac{ dx }{ \sqrt{1-36x^2} }\]

OpenStudy (anonymous):

ok so you can rewrite that as \[\int\limits\limits_{}^{}\frac{ dx }{ \sqrt{36(\frac{ 1 }{ 36 }-x^{2}}) }=\frac{ 1 }{ 6 } \int\limits\limits_{}^{}\frac{ dx }{ \sqrt{(\frac{ 1 }{ 6 })^{2}-x^{2}} }\] now using table integral \[\int\limits_{}^{}\frac{ dx }{ \sqrt{a^{2}-x^{2}} }=\sin^{-1}(\frac{ x }{ a })\] yu can solve it :)

OpenStudy (bluesea24):

yes solved it ty :)